OpenStudy (anonymous):

Solve 2(3 – 2x) = x – 9. Show all work

OpenStudy (anonymous):

What's the first step?

OpenStudy (anonymous):

getting rid of parenthesis?

OpenStudy (anonymous):

Yep! How do you do that?

OpenStudy (anonymous):

distributive property?

OpenStudy (anonymous):

distributive

OpenStudy (anonymous):

2(3 – 2x) = x – 9. 6 - 4x = x -9 15 = 5x x = 15/5 =3

OpenStudy (anonymous):

Alright, so @itsmichi95 what do you get from the distribution? @Yahoo!, please let itsmichi95 solve this so she can figure out how it works.